### v1

# Searching for new physics with 136Xe double beta decay spectrum in PandaX-4T

Searching for new physics with 136Xe double beta decay spectrum in PandaX-4T
Zhe YuanKey Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Zihao BoState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Wei ChenState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xun ChenState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yunhua Chen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Chen ChengSchool of Physics, Beihang University, Beijing 102206, China
Xiangyi Cui
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Manna Deng
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Yingjie Fan
Department of Physics, Yantai University, Yantai 264005, China
Deqing Fang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Xuanye FuState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Zhixing Gao
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Yujie GeSino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Lisheng GengSchool of Physics, Beihang University, Beijing 102206, China
Peng Huanwu Collaborative Center for Research and Education, Beihang University, Beijing 100191, China
International Research Center for Nuclei and Particles in the Cosmos & Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 100191, China
Southern Center for Nuclear-Science Theory (SCNT), Institute of Modern Physics, Chinese Academy of Sciences, Huizhou 516000, China
Karl Giboni
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xunan GuoSchool of Physics, Beihang University, Beijing 102206, China
Xuyuan Guo
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Zichao Guo
School of Physics, Beihang University, Beijing 102206, China
Chencheng HanState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Ke Hanke.han@sjtu.edu.cnState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Changda He
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinrong He
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Houqi Huang
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Junting HuangState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yule Huang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Ruquan Hou
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xiangdong JiDepartment of Physics, University of Maryland, College Park, Maryland 20742, USA
Yonglin JuSchool of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xiaorun Lan
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Chenxiang Li
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jiafu Li
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Mingchuan Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Peiyuan LiState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Shuaijie LiYalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Tao LiSJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Yangdong Li
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Zhiyuan Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Qing LinState Key Laboratory of Particle Detection and Electronics, University of Science and Technology of China, Hefei 230026, China
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Jianglai Liujianglai.liu@sjtu.edu.cnState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yuanchun Liu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Congcong Lu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiaoying Lu
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Lingyin Luo
School of Physics, Peking University, Beijing 100871, China
Yunyang Luo
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Yugang MaKey Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Yajun Mao
School of Physics, Peking University, Beijing 100871, China
Yue MengState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Binyu Pang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Ningchun Qi
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Zhicheng Qian
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xiangxiang Ren
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Dong Shan
School of Physics, Nankai University, Tianjin 300071, China
Xiaofeng Shang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xiyuan ShaoSchool of Physics, Nankai University, Tianjin 300071, China
Guofang Shen
School of Physics, Beihang University, Beijing 102206, China
Manbin Shen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Wenliang Sun
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xuyan SunState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Yi TaoSchool of Science, Sun Yat-Sen University, Shenzhen 518107, China
Yueqiang Tian
School of Physics, Beihang University, Beijing 102206, China
Yuxin Tian
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Anqing Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Guanbo Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Hao WangState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Haoyu WangState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jiamin Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Lei Wang
College of Nuclear Technology and Automation Engineering, Chengdu University of Technology, Chengdu 610059, China
Meng WangResearch Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Qiuhong WangKey Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Shaobo WangState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Shibo Wang
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Siguang Wang
School of Physics, Peking University, Beijing 100871, China
Wei WangSino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Xu Wang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Zhou WangState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yuehuan WeiSino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai 519082, China
Weihao Wu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Yuan Wu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Mengjiao XiaoState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xiang Xiaoxiaox93@mail.sysu.edu.cnSchool of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Kaizhi Xiong
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Jianqin Xu
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Yifan Xu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Shunyu Yao
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Binbin YanState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Xiyu Yan
School of Physics and Astronomy, Sun Yat-Sen University, Zhuhai 519082, China
Yong Yang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Peihua Ye
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Chunxu Yu
School of Physics, Nankai University, Tianjin 300071, China
Ying Yuan
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Youhui Yun
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Xinning Zeng
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Minzhen Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Peng Zhang
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Shibo ZhangState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Siyuan Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Shu Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Tao Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Wei Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Yang Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Yingxin Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, China
Yuanyuan Zhang
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Li ZhaoState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Kangkang Zhao
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Jifang Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Jiaxu Zhou
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai 200240, China
Jiayi Zhou
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Ning ZhouState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
Xiaopeng ZhouSchool of Physics, Beihang University, Beijing 102206, China
Zhizhen Zhou
State Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
Chenhui Zhu
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Dong-Liang FangInstitute of Modern Physics, Chinese Academy of Sciences, Lanzhou 730000, China
School of Nuclear Science and Technology, University of Chinese Academy of Sciences, Beijing 100049, China
Yu-Feng LiInstitute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China
School of Physical Sciences, University of Chinese Academy of Sciences, Beijing 100049, China
(December 4, 2025)
Abstract
The continuous spectrum of double beta decay () provides a sensitive probe to test the predictions of the Standard Model and to search for signatures of new physics beyond it.
We present a comprehensive analysis of the 136Xe spectrum utilizing kgyr of 136Xe exposure from the PandaX-4T experiment.
The analysis yields the most precise measurement to date of the 136Xe two-neutrino double beta decay () half-life, years, the uncertainty of which is reduced by a factor of two compared to our previous result.
We measure the parameter , defined as the ratio between the subleading and leading components of the 136Xe nuclear matrix element, to be , which is consistent with theoretical predictions.
We also search for Majoron-emitting modes of 136Xe , establishing the most stringent limit for the spectral index .
Double beta decay () is a rare nuclear process [DBD:origin].
If two neutrinos are emitted during this process (), it is consistent with the electroweak interaction framework of the Standard Model of particle physics.
In contrast, if no neutrinos are emitted, the process is referred to as neutrinoless double beta decay (), which can only occur if neutrinos are Majorana particles [NLDBD:probability, NLDBD:review1, NLDBD:review2].
Additionally, if one or two bosons are emitted in processes, they are typically referred to as “Majorons” [Majoron:goldstone, Majoron:Left-handed, Majoron:Unconventional].
Although has been experimentally observed and measured in several studies, such as those involving 136Xe [EXO-200:Xe136dbd2014, KamLAND-Zen:kesi31, NEXT:Xe136dbd, PandaX:2022kwg], neither nor Majoron-emitting has been detected.
Recent experiments have established the most stringent lower limits on the half-life of different isotopes, such as 136Xe [KamLAND:Xenldbd2023, EXO-200:2019nldbd], 76Ge [GERDA:Genldbd2020, Majorana:2023Genldbd, CDEX:2024Genldbd], and 130Te [CUORE:2022Tenldbd].
The half-life is linked to the effective Majorana neutrino mass by [NLDBD:review2]
where is the electron mass, the axial-vector coupling constant, the nuclear matrix element (NME), and the phase space factor.
The current understanding of NME remains incomplete, with significant discrepancies observed in different nuclear many-body approaches and potential influences of quenching effects [NME:Status, Xe136nldbd:2015review].
Since and share the same initial and final nuclear states, accurately determining NME is essential to constrain NME predictions [kesi31:theory].
A refined expression for the half-life is presented as follows [kesi31:theory]:
where is the effective axial-vector coupling constant.
In contrast to the standard phase space factor , the phase space factor accounts for the dependence on lepton energies from the energy denominators of NMEs.
The parameter is defined as , where the subleading NME is only sensitive to contributions from the lightest states of the intermediate nucleus, while the NME is also sensitive to higher-lying states.
The experimentally accessible parameter , which can be measured through electron energy spectrum fits, allows for the discrimination between different theoretical models and thus provides new insights into NME calculations.
Majorons are potential dark matter candidates and play roles in various cosmological and astrophysical processes [Majoron:darkmatter1, Majoron:darkmatter2, Majoron:darkmatter3].
The Majoron-emitting encompasses a range of modes, which are distinguished by the number of emitted bosons, the leptonic charge, whether lepton number is violated, and whether the Majoron is a Goldstone boson.
Different modes may produce identical summed electron energy spectra, which are classified into four types based on the spectral index 1, 2, 3 and 7 [Majoron:n137, Majoron:n2].
Previous experimental searches have established stringent constraints on these processes with 136Xe [KamLAND-Zen:Majoron, EXO-200:Majoron2021].
Previous searches for and Majoron-emitting of 136Xe have been limited to relatively high energy thresholds, specifically between 500 keV and 800 keV [KamLAND-Zen:kesi31, KamLAND-Zen:Majoron, EXO-200:Majoron2021].
Consequently, the lower energy region remains completely unexplored.
The PandaX-4T detector, which uses a natural xenon target and was initially designed for dark matter searches in the low-energy region, has expanded its detection capability to the MeV range [PandaX:2022kwg, PandaX-4T:Xe136nldbd].
In this paper, we leverage the nearly complete spectrum to perform the measurement of and the searches for Majoron-emitting of 136Xe, using a dataset of the commissioning run (Run0) and the first science run (Run1) of the PandaX-4T experiment [PandaX:2025WIMP, PandaX:SuperWIMP, PandaX-4T:Xe136nldbd] with a total 136Xe exposure of kgyr.
The energy region of interest (ROI) for the spectrum fit is from 20 keV to 2800 keV.
The PandaX-4T detector is a dual-phase time projection chamber (TPC) containing 3.7 tonnes of natural xenon in its active volume.
The TPC has a vertical height of 1.185 m and is enclosed by a field cage along its lateral surfaces.
The lateral surfaces of the TPC are lined with reflective Teflon panels, arranged in a regular 24-sided polygon with an inscribed diameter of 1.185 m.
Three electrode grids (cathode, gate, and anode) are vertically stacked within the TPC to generate a drift field and an extraction field.
Two arrays of 3-inch Hamamatsu photomultiplier tubes (PMTs) are placed on the top and bottom as photosensors.
More detailed description of the detector can be found in Ref. [PandaX-4T:2021bab].
Energy deposition in liquid xenon (LXe) produces a prompt scintillation signal () and ionization electrons.
These electrons drift upward into the gaseous xenon phase, where they generate a delayed electroluminescence signal () via proportional scintillation.
Energy reconstruction utilizes the area of and .
Spatial reconstruction determines the vertical (z axis) position via drift velocity and the time delay between the and signals, while the horizontal (x-y axis) coordinates are obtained through maximum likelihood estimation of the charge distribution pattern of observed in the top PMT array.
The data production and event selection procedures follow those described in Refs. [PandaX-4T:Xe136nldbd, PandaX:SuperWIMP, PandaX:2024med, PandaX:2023ggs].
The fiducial volume (FV) from the previous study [PandaX:SuperWIMP], defined as the innermost cleanest part of the detector, is adopted in this analysis.
The fiducial mass is determined to be kg for Run0 and kg for Run1, by scaling the ratio of events in the FV and in the whole TPC.
The uncertainty arises from the LXe density and the difference between the geometrically-calculated volume and the volume scaled by the [PandaX:SuperWIMP].
The potential shift between the reconstructed energy and the true energy as well as the energy resolution are modelled using the 164 keV peak (from Xe), along with the 236 keV peaks (from Xe and Xe), the 1460 keV peak (from 40K), and the 2615 keV peak (from 232Th), obtained from data in the control region outside of the FV.
This ensures that the energy response model is completely uncorrelated with the data used in the final fit.
Five parameters are used to model the energy response, independently for Run0 and Run1.
The energy resolution is modeled as a Gaussian function, with the width given by , where is the reconstructed energy in keV.
The residual energy nonlinearity is modeled as , where is the true energy.
The calibrated values and uncertainties (Table 1), along with the covariance matrix , are used to fit the Run0 and Run1 data.
The specific approach is consistent with that in Ref. [PandaX:SuperWIMP].
Table 1:
Summary of sources of systematic uncertainties.
denotes the 5-parameter detector response model (see text), with its means and uncertainties determined from mono-energetic peaks obtained from the peripheral data of FV.
The total detector efficiency comprises three components: the single-site (SS) fraction, quality cut (QC) efficiency, and ROI acceptance, estimated with methods inherited from Ref. [PandaX:SuperWIMP] and summarized in Table 1.
The SS fraction is defined as the ratio of SS to the sum of SS and multi-site (MS) energy deposition events in the detector.
We validate the SS fraction within the ROI using 232Th calibration data together with 232Th events simulated with the BambooMC simulation framework [Chen:2021asx].
In this analysis, the QC efficiency of Run0 and Run1 are ()% and ()% respectively.
To generate the signal spectra for the fit, the theoretical spectra of and in Eq. 2 are calculated from the numerical wave functions which are the solutions of Dirac equation [NLDBDtheory:leftright], obtained with the package RADIAL [Diracwave] by assuming a uniform distribution of the nuclear charge inside the nucleus.
The detector-response-convolved spectra are then simulated by the BambooMC package.
For and Majoron-emitting spectra, we use the Decay0 package [BxDecay0] to generate events with momentum of two electrons, and then simulate the detector response with BambooMC package as well.
Figure 1: Spectra for , , and the Majoron-emitting modes with = 1, 2, 3, and 7, respectively. Detector response model is taken into account.
The obtained spectra with detector response incorporated are shown in Fig. 1.
The ROI acceptance for 136Xe in the fit varies with , while the acceptances for Majoron-emitting modes are 100.00% (), 100.00% (), 99.99% (), and 99.95% ().
Background events originate from three primary sources: liquid xenon, external structure, and solar neutrinos, as summarized in Table 2.
124Xe and 125I are taken from the measurements reported in Ref. [PandaX:2024xpq].
For xenon isotopes in Run0, including 127Xe, 129mXe and 131mXe, we adopt the xenon-evolution analysis of Ref. [PandaX:SuperWIMP].
All remaining xenon isotopes are left free in the fit.
85Kr concentration is determined via - cascade tagging through the isomeric 85mRb [Collon:2004xs].
214Pb and 212Pb components are also treated as free parameters.
The external structure, including both the detector materials and the stainless steel platform (SSP), contributes to the background.
Activities of 232Th, 238U, 60Co, and 40K in the detector materials are taken from measurements using high-purity germanium (HPGe) counting stations [PandaX:materialHGe], while activities of 232Th and 238U in SSP are taken from Ref. [PandaX:SSP].
The solar and 7Be neutrinos contributions are taken from Refs. [Chen:2016eab, BOREXINO:2014pcl].
Table 2: The background contributions in the ROI for Run0 and Run1.
The fitted counts are obtained from the 136Xe half-life measurement, the measurement and the Majoron-emitting search with mode .
Components
Expected
Majoron
136Xe
SSP 232Th
SSP 238U
60Co
40K
232Th
238U
164 keV (Run0)
208 keV (Run0)
236 keV (Run0)
380 keV (Run0)
408 keV (Run0)
125I(Run0)
125Xe (Run0)
float
214Pb (Run0)
float
133Xe (Run0)
float
212Pb (Run0)
float
85Kr (Run0)
164 keV (Run1)
float
236 keV (Run1)
float
125I (Run1)
214Pb (Run1)
float
212Pb (Run1)
float
85Kr (Run1)
124Xe
Be
Assuming no new physics is observed in this study, a one-dimensional binned likelihood fit to measure the half-life of 136Xe is performed first, as shown in Fig. 2.
The likelihood is constructed as
where and are the expected and observed numbers of events in the energy bin for Run-, respectively.
in Run- is defined as
where and are the counts of signal and background component , respectively.
The corresponding and are the bin values of the normalized energy spectrum convolved with the five-parameter energy response model.
The Gaussian penalty term of the energy response contains the five-parameter and the covariant matrix in Run-.
The Gaussian penalty terms are used to constrain the nuisance parameters , , and , which represent the relative uncertainties in the overall efficiency, the signal selection, and background model (see Table 2), respectively.
The radioactivities of external structures, 124Xe concentrations, and solar and 7Be fluxes are identical in Run0 and Run1.
Other background components are treated independently for Run0 and Run1.
Figure 2: The SS data spectra and the fit for the 136Xe half-life are shown for Run0 (top) and Run1 (bottom) from 20 keV to 2800 keV with a bin size of 4 keV. The horizontal axis represents the reconstructed energy in the data. Xe∗ includes the contributions from 124Xe, 125Xe, 127Xe, 129mXe, 131mXe, and 133Xe. The lower panel shows the residuals together with , and bands.
The observed backgrounds are in general agreement with expectations, as shown in Table 2.
The fitted contributions from Kr in Run1 and SSP 232Th are pulled by 1.8 and 1.2, respectively, relative to their input values.
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
The half-life of 136Xe is measured as years, representing the most precise result so far, with a total uncertainty smaller than that of any previous measurement.
It is worth noting that this result is slightly smaller than our previous result from Run0 data [PandaX:2022kwg], due to the use of the updated and slightly smaller 136Xe abundance of [PandaX:SSP], compared to the previous value of .
The 136Xe spectrum is treated as the background in the search for Majoron-emitting signals, as shown later.
Figure 3:
Two-dimensional profile likelihood scan for 136Xe rate and .
The black dot indicates the best-fit value of .
The contours correspond to the 68.3, 90, 95.4, and 99.7 CL.
In the fit of , we allow both the half-life of 136Xe and (i.e., the shape of 136Xe ) to float freely.
The likelihood is almost the same as Eq. 3, but with a different definition of in Run- which is modified as
where corresponds to the PDF corresponding to the first term of the Taylor expansion of the decay rate of , and corresponds to the PDF corresponding to the second term.
The parameter , which comprises , and in Eq. 2 as well as the total detection efficiencies, varies linearly with .
The rest of the parameters remain the same as in Eq. 4.
The backgrounds, also summarized in Table 2, are consistent with expectations, except for the contributions from Kr in Run1 and SSP 232Th, which have pulls of 1.6 and 1.2, respectively.
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
We obtained a best fit of , consistent with theoretical predictions [kesi31:theory] and the KamLAND-Zen results [KamLAND-Zen:kesi31].
Fig. 3 shows the two-dimensional confidence intervals for the 136Xe rate and , revealing only a slight positive correlation.
The inclusion of the second-order contribution is found to have a negligible impact on the 136Xe decay rate.
The 136Xe decay half-life is estimated to be years.
This result is consistent with the 136Xe -only fit, when the second-order contribution was not considered.
Figure 4: Spectra fits with Majoron-emitting decay signals ().
The figure is the same as Fig. 2 but with the upper limits (90% CL) of signals illustrated as hatched histograms. The lower panel shows the residuals together with , and bands.
In the fit of Majoron-emitting , the likelihood is similar to Eq. 3, except that 136Xe is moved from the signal term to the background term, and then Majoron-emitting is introduced as the signal.
As an example, the best-fit spectrum for mode is shown in Fig. 4.
No statistically significant evidence for Majoron-emitting decays is observed for any mode considered here, and the lower limits on Majoron-emitting half-lives are derived at the confidence level (CL), as summarized in Table 3.
The extension of the lower bound of the ROI from 600 keV (as in EXO-200 [EXO-200:Majoron2021]) to 20 keV considerably improves the sensitivity to mode , where the Majoron-emitting spectrum peaks at approximately 600 keV, leading to the most stringent constraint.
However, our sensitivity to modes 1, 2, and 3 is limited due to our 136Xe exposure being only 0.16 times that of EXO-200 [EXO-200:Majoron2021] and our higher background level.
Similarly, the backgrounds are consistent with expectations, with only the contributions from Kr in Run1 and SSP 232Th pulled by 1.7 and 1.2, respectively (Table 2).
The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
Table 3: The 90% CL lower limits on half-lives for different Majoron-emitting decay models and the comparison to EXO-200 results [EXO-200:Majoron2021]
Decay mode
This work (yr)
EXO-200 (2021) (yr)
n = 1
n = 2
n = 3
n = 7
In summary, we perform a measurement of in and a search for Majoron-emitting for 136Xe with a total 136Xe exposure of kgyr using PandaX-4T data.
The most precise measurement of the 136Xe half-life to date is achieved thanks to the nearly complete spectrum used in the fit.
Our measurement of is consistent with theoretical predictions and the results from KamLAND-Zen.
The most competitive results for the Majoron-emitting mode with are obtained.
Following detector upgrades in 2023, PandaX-4T has been taking data in the second science run (Run2).
The capability to reconstruct the nearly complete 136Xe spectrum will allow us to fully exploit the PandaX-4T dataset, significantly enhancing its physics potential.
This project is supported in part by grants from National Key R&D Program of China (Nos. 2023YFA1606200, 2023YFA1606202), National Science Foundation of China (Nos. 12090060, 12090062, 12305121, U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 21TQ1400218, 22JC1410100, 23JC1410200, ZJ2023-ZD-003). We thank for the support by the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.

### v2

@@
-Searching for new physics with 136Xe double beta decay spectrum in PandaX-4T
-Searching for new physics with 136Xe double beta decay spectrum in PandaX-4T
+Precise 136Xe Double Beta Decay Measurement in PandaX-4T with Implications on the Nuclear Matrix Elements and Majorons
+Precise 136Xe Double Beta Decay Measurement in PandaX-4T with Implications on the Nuclear Matrix Elements and Majorons
 Zhe YuanKey Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
@@
 Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province, Liangshan 615000, China
-Xiangdong JiDepartment of Physics, University of Maryland, College Park, Maryland 20742, USA
+Xiangdong JiState Key Laboratory of Dark Matter Physics, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Tsung-Dao Lee Institute & School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 201210, China
+Department of Physics, University of Maryland, College Park, Maryland 20742, USA
 Yonglin JuSchool of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
@@
 School of Physical Sciences, University of Chinese Academy of Sciences, Beijing 100049, China
-(December 4, 2025)
 Abstract
 The continuous spectrum of double beta decay () provides a sensitive probe to test the predictions of the standard model and to search for signatures of new physics beyond it.
-We present a comprehensive analysis of the 136Xe spectrum utilizing kgyr of 136Xe exposure from the PandaX-4T experiment.
-The analysis yields the most precise measurement to date of the 136Xe two-neutrino double beta decay () half-life, years, the uncertainty of which is reduced by a factor of two compared to our previous result.
+We present a comprehensive analysis of the 136Xe spectrum utilizing of 136Xe exposure from the PandaX-4T experiment.
+The analysis yields the most precise measurement to date of the 136Xe two-neutrino double beta decay () half-life, years, the uncertainty of which is reduced by a factor of 2 compared to our previous result.
 We measure the parameter , defined as the ratio between the subleading and leading components of the 136Xe nuclear matrix element, to be , which is consistent with theoretical predictions.
 We also search for Majoron-emitting modes of 136Xe , establishing the most stringent limit for the spectral index .
-Double beta decay () is a rare nuclear process [DBD:origin].
+Double beta decay () is a rare nuclear process [25].
 If two neutrinos are emitted during this process (), it is consistent with the electroweak interaction framework of the standard model of particle physics.
-In contrast, if no neutrinos are emitted, the process is referred to as neutrinoless double beta decay (), which can only occur if neutrinos are Majorana particles [NLDBD:probability, NLDBD:review1, NLDBD:review2].
-Additionally, if one or two bosons are emitted in processes, they are typically referred to as “Majorons” [Majoron:goldstone, Majoron:Left-handed, Majoron:Unconventional].
-Although has been experimentally observed and measured in several studies, such as those involving 136Xe [EXO-200:Xe136dbd2014, KamLAND-Zen:kesi31, NEXT:Xe136dbd, PandaX:2022kwg], neither nor Majoron-emitting has been detected.
-Recent experiments have established the most stringent lower limits on the half-life of different isotopes, such as 136Xe [KamLAND:Xenldbd2023, EXO-200:2019nldbd], 76Ge [GERDA:Genldbd2020, Majorana:2023Genldbd, CDEX:2024Genldbd], and 130Te [CUORE:2022Tenldbd].
-The half-life is linked to the effective Majorana neutrino mass by [NLDBD:review2]
+In contrast, if no neutrinos are emitted, the process is referred to as neutrinoless double beta decay (), which can occur only if neutrinos are Majorana particles [20, 39, 5, 26].
+Additionally, if one or two bosons are emitted in processes, they are typically referred to as “Majorons” [16, 23, 24].
+Although has been experimentally observed and measured in several studies, such as those involving 136Xe [7, 22, 35, 40], neither nor Majoron-emitting has been detected.
+Recent experiments have established the most stringent lower limits on the half-life of different isotopes, such as 136Xe [1, 8], 76Ge [4, 9, 44, 2], and 130Te [3].
+The half-life is linked to the effective Majorana neutrino mass by [26]
 where is the electron mass, the axial-vector coupling constant, the nuclear matrix element (NME), and the phase space factor.
-The current understanding of NME remains incomplete, with significant discrepancies observed in different nuclear many-body approaches and potential influences of quenching effects [NME:Status, Xe136nldbd:2015review].
-Since and share the same initial and final nuclear states, accurately determining NME is essential to constrain NME predictions [kesi31:theory].
-A refined expression for the half-life is presented as follows [kesi31:theory]:
-where is the effective axial-vector coupling constant.
-In contrast to the standard phase space factor , the phase space factor accounts for the dependence on lepton energies from the energy denominators of NMEs.
-The parameter is defined as , where the subleading NME is only sensitive to contributions from the lightest states of the intermediate nucleus, while the NME is also sensitive to higher-lying states.
-The experimentally accessible parameter , which can be measured through electron energy spectrum fits, allows for the discrimination between different theoretical models and thus provides new insights into NME calculations.
-Majorons are potential dark matter candidates and play roles in various cosmological and astrophysical processes [Majoron:darkmatter1, Majoron:darkmatter2, Majoron:darkmatter3].
+The current understanding of NME remains incomplete, with significant discrepancies observed in different nuclear many-body approaches and potential influences of quenching effects [19, 18].
+Since and share the same initial and final nuclear states, accurately determining NME is essential to constrain NME predictions [41].
+A refined expression for the half-life, retaining terms up to second order in the Taylor expansion, is presented as follows [41]:
+where is the effective axial-vector coupling constant and denotes the leading-order Gamow-Teller component of the NME.
+In contrast to the standard phase space factor , the phase space factor accounts for the dependence on lepton energies of the energy denominators of NMEs.
+The parameter is defined as , where the subleading NME is sensitive only to contributions from the lightest states of the intermediate nucleus, while the NME is also sensitive to higher-lying states.
+The experimentally accessible parameter , which can be measured through electron energy spectrum fits, allows for the discrimination between different theoretical models, and thus, provides new insights into NME calculations.
+Specifically, a zero (nonzero) value of implies the higher (single) state dominance hypothesis of [41].
+Majorons are potential dark matter candidates and play roles in various cosmological and astrophysical processes [28, 10, 29].
 The Majoron-emitting encompasses a range of modes, which are distinguished by the number of emitted bosons, the leptonic charge, whether lepton number is violated, and whether the Majoron is a Goldstone boson.
-Different modes may produce identical summed electron energy spectra, which are classified into four types based on the spectral index 1, 2, 3 and 7 [Majoron:n137, Majoron:n2].
-Previous experimental searches have established stringent constraints on these processes with 136Xe [KamLAND-Zen:Majoron, EXO-200:Majoron2021].
-Previous searches for and Majoron-emitting of 136Xe have been limited to relatively high energy thresholds, specifically between 500 keV and 800 keV [KamLAND-Zen:kesi31, KamLAND-Zen:Majoron, EXO-200:Majoron2021].
+Different modes may produce identical summed electron energy spectra, which are classified into four types based on the spectral index 1, 2, 3, and 7 [27, 34].
+Previous experimental searches have established stringent constraints on these processes with 136Xe [21, 6].
+Previous measurements of and searches for Majoron-emitting of 136Xe have been limited to relatively high energy thresholds, specifically between 500 keV and 800 keV [22, 21, 6].
 Consequently, the lower energy region remains completely unexplored.
-The PandaX-4T detector, which uses a natural xenon target and was initially designed for dark matter searches in the low-energy region, has expanded its detection capability to the MeV range [PandaX:2022kwg, PandaX-4T:Xe136nldbd].
-In this paper, we leverage the nearly complete spectrum to perform the measurement of and the searches for Majoron-emitting of 136Xe, using a dataset of the commissioning run (Run0) and the first science run (Run1) of the PandaX-4T experiment [PandaX:2025WIMP, PandaX:SuperWIMP, PandaX-4T:Xe136nldbd] with a total 136Xe exposure of kgyr.
-The energy region of interest (ROI) for the spectrum fit is from 20 keV to 2800 keV.
-The PandaX-4T detector is a dual-phase time projection chamber (TPC) containing 3.7 tonnes of natural xenon in its active volume.
+The PandaX-4T detector, which uses a natural xenon target and was initially designed for dark matter searches in the low-energy region, has expanded its detection capability to the MeV range [40, 45].
+In this Letter, we leverage the nearly complete spectrum to perform the measurement of and the searches for Majoron-emitting modes of of 136Xe, using a dataset of the commissioning run (Run0, 94.8 days of live time) and the first science run (Run1, 163.5 days) of the PandaX-4T experiment [12, 30, 45] with a total 136Xe exposure of .
+The energy region of interest (ROI) for the spectrum fit is from 20 keV to
+2800 keV.
+The PandaX-4T detector is a dual-phase time projection chamber (TPC) containing 3.7 tons of natural xenon in its active volume.
 The TPC has a vertical height of 1.185 m and is enclosed by a field cage along its lateral surfaces.
@@
 Two arrays of 3-inch Hamamatsu photomultiplier tubes (PMTs) are placed on the top and bottom as photosensors.
-More detailed description of the detector can be found in Ref. [PandaX-4T:2021bab].
+A more detailed description of the detector can be found in Ref. [33].
 Energy deposition in liquid xenon (LXe) produces a prompt scintillation signal () and ionization electrons.
 These electrons drift upward into the gaseous xenon phase, where they generate a delayed electroluminescence signal () via proportional scintillation.
-Energy reconstruction utilizes the area of and .
-Spatial reconstruction determines the vertical (z axis) position via drift velocity and the time delay between the and signals, while the horizontal (x-y axis) coordinates are obtained through maximum likelihood estimation of the charge distribution pattern of observed in the top PMT array.
-The data production and event selection procedures follow those described in Refs. [PandaX-4T:Xe136nldbd, PandaX:SuperWIMP, PandaX:2024med, PandaX:2023ggs].
-The fiducial volume (FV) from the previous study [PandaX:SuperWIMP], defined as the innermost cleanest part of the detector, is adopted in this analysis.
+Energy reconstruction utilizes the charge of and , where stands for detected by the bottom PMT array, following the procedure detailed in Ref. [45, 30, 32, 43].
+Spatial reconstruction determines the vertical (-axis) position via drift velocity and the time delay between the and signals, while the horizontal (- and -axes) coordinates are obtained through maximum likelihood estimation of the charge distribution pattern of observed in the top PMT array [30].
+The data production and event selection procedures follow those described in Refs. [45, 30, 32, 43].
+The fiducial volume (FV) from the previous study [30], defined as the innermost, cleanest part of the detector, is adopted in this analysis.
 The fiducial mass is determined to be kg for Run0 and kg for Run1, by scaling the ratio of events in the FV and in the whole TPC.
-The uncertainty arises from the LXe density and the difference between the geometrically-calculated volume and the volume scaled by the [PandaX:SuperWIMP].
-The potential shift between the reconstructed energy and the true energy as well as the energy resolution are modelled using the 164 keV peak (from Xe), along with the 236 keV peaks (from Xe and Xe), the 1460 keV peak (from 40K), and the 2615 keV peak (from 232Th), obtained from data in the control region outside of the FV.
-This ensures that the energy response model is completely uncorrelated with the data used in the final fit.
-Five parameters are used to model the energy response, independently for Run0 and Run1.
-The energy resolution is modeled as a Gaussian function, with the width given by , where is the reconstructed energy in keV.
-The residual energy nonlinearity is modeled as , where is the true energy.
-The calibrated values and uncertainties (Table 1), along with the covariance matrix , are used to fit the Run0 and Run1 data.
-The specific approach is consistent with that in Ref. [PandaX:SuperWIMP].
+The uncertainty arises from the LXe density and the difference between the geometrically calculated volume and the volume scaled by the uniformly distributed calibration source [30].
+Compared with our previous study of the 136Xe decay [40], the present analysis extends the analysis window down to 20 keV.
+Similar to Ref.
+[30], monoenergetic peaks at 41.5 keV (from ), 164 keV (Xe), and 236 keV (127Xe and Xe) are used for energy calibration, together with high-energy peaks from 137Cs, 60Co, 40K, and the 232Th decay chain.
+The energy response is modeled using five parameters, which account for a possible shift between the reconstructed and true energies as well as the energy resolution [30].
+The energy resolution is parametrized by a Gaussian function with a width given by
+, where is the reconstructed energy in keV. The residual energy nonlinearity is modeled as , where denotes the true energy.
+The calibrated parameter values and their uncertainties, (Table 1), together with the corresponding covariance matrix , are used to fit the Run0 and Run1 data independently.
+Due to the lack of sufficient high-energy calibration data within the fiducial volume, the energy response model is derived from physics data in a control region outside the FV, providing a closely matched but statistically uncorrelated dataset for the final fit [45].
 Table 1:
 Summary of sources of systematic uncertainties.
-denotes the 5-parameter detector response model (see text), with its means and uncertainties determined from mono-energetic peaks obtained from the peripheral data of FV.
-The total detector efficiency comprises three components: the single-site (SS) fraction, quality cut (QC) efficiency, and ROI acceptance, estimated with methods inherited from Ref.
-[PandaX:SuperWIMP] and summarized in Table 1.
-The SS fraction is defined as the ratio of SS to the sum of SS and multi-site (MS) energy deposition events in the detector.
-We validate the SS fraction within the ROI using 232Th calibration data together with 232Th events simulated with the BambooMC simulation framework [Chen:2021asx].
-In this analysis, the QC efficiency of Run0 and Run1 are ()% and ()% respectively.
-To generate the signal spectra for the fit, the theoretical spectra of and in Eq.
-2 are calculated from the numerical wave functions which are the solutions of Dirac equation [NLDBDtheory:leftright], obtained with the package RADIAL [Diracwave] by assuming a uniform distribution of the nuclear charge inside the nucleus.
+denotes the 5-parameter detector response model (see text), with its means and uncertainties determined from monoenergetic peaks obtained from the control data outside of the FV.
+Sources
+Run0 (94.8 d)
+Run1 (163.5 d)
+Detectorresponse
+[keV-1]
+[keV]
+Overallefficiency
+SS fraction ()
+Quality cut
+Signalselection
+LXe density [g/cm3]
+136Xe abundance
+FV mass [kg]
+The total detection efficiency comprises three components: the single-site (SS) fraction, quality cut (QC) efficiency, and ROI acceptance, estimated with methods inherited from Ref.
+[30] and summarized in Table 1.
+The SS fraction is defined as the ratio of SS to the sum of SS and multisite (MS) energy deposition events in the detector.
+We validate the SS fraction within the ROI using 232Th calibration data together with 232Th events simulated with the BambooMC simulation framework [15].
+In this analysis, the QC efficiency of Run0 and Run1 are ()% and ()%, respectively.
+To generate the signal spectra for the fit, the theoretical spectra of the first and second terms of the Taylor expansion in Eq.
+(2) are calculated from the numerical wave functions, which are the solutions of the Dirac equation [42], obtained with the package RADIAL [38] by assuming a uniform distribution of the nuclear charge inside the nucleus.
 The detector-response-convolved spectra are then simulated by the BambooMC package.
-For and Majoron-emitting spectra, we use the Decay0 package [BxDecay0] to generate events with momentum of two electrons, and then simulate the detector response with BambooMC package as well.
+For and Majoron-emitting spectra, we use the Decay0 package [36] to generate events with momenta of two electrons, and then, simulate the detector response with the BambooMC package as well.
 Figure 1: Spectra for , , and the Majoron-emitting modes with = 1, 2, 3, and 7, respectively. Detector response model is taken into account.
@@
 The ROI acceptance for 136Xe in the fit varies with , while the acceptances for Majoron-emitting modes are 100.00% (), 100.00% (), 99.99% (), and 99.95% ().
-Background events originate from three primary sources: liquid xenon, external structure, and solar neutrinos, as summarized in Table 2.
-124Xe and 125I are taken from the measurements reported in Ref. [PandaX:2024xpq].
-For xenon isotopes in Run0, including 127Xe, 129mXe and 131mXe, we adopt the xenon-evolution analysis of Ref. [PandaX:SuperWIMP].
-All remaining xenon isotopes are left free in the fit.
-85Kr concentration is determined via - cascade tagging through the isomeric 85mRb [Collon:2004xs].
+Background events originate from three primary sources: liquid xenon, the detector material and the stainless steel platform (SSP) material, and solar neutrinos, as summarized in Table 2.
+124Xe and 125I are taken from the measurements reported in Ref. [13].
+For 127Xe, 129mXe, and 131mXe in Run0, we adopt the xenon-evolution analysis from Ref. [30].
+In Run1, due to the absence of neutron calibration, the activities of these xenon isotopes are very low, and their time evolution cannot be determined precisely.
+Therefore, we allow their activities to float in the fit.
+The activities of 125Xe and 133Xe are left free in Run0 and are negligible in Run1.
+85Kr concentration is determined via - cascade tagging through the isomeric 85mRb [17].
 214Pb and 212Pb components are also treated as free parameters.
-The external structure, including both the detector materials and the stainless steel platform (SSP), contributes to the background.
-Activities of 232Th, 238U, 60Co, and 40K in the detector materials are taken from measurements using high-purity germanium (HPGe) counting stations [PandaX:materialHGe], while activities of 232Th and 238U in SSP are taken from Ref.
-[PandaX:SSP].
-The solar and 7Be neutrinos contributions are taken from Refs. [Chen:2016eab, BOREXINO:2014pcl].
+The material of the detector and the SSP contributes to the background.
+Activities of 232Th, 238U, 60Co, and 40K in the detector components are taken from measurements using high-purity germanium counting stations [37], while activities of 232Th and 238U in SSP components are taken from Ref.
+[31].
+The solar and 7Be neutrino contributions are taken from Refs. [14, 11].
 Table 2: The background contributions in the ROI for Run0 and Run1.
-The fitted counts are obtained from the 136Xe half-life measurement, the measurement and the Majoron-emitting search with mode .
+The fitted counts are obtained from the 136Xe half-life measurement, the measurement, and the Majoron-emitting search with mode .
 Components
@@
 SSP 238U
-60Co
-40K
-232Th
-238U
+Detector 60Co
+Detector 40K
+Detector 232Th
+Detector 238U
 164 keV (Run0)
@@
 Be
-Assuming no new physics is observed in this study, a one-dimensional binned likelihood fit to measure the half-life of 136Xe is performed first, as shown in Fig. 2.
+Assuming no new physics is observed in this study, a one-dimensional binned likelihood fit is first performed to measure the half-life of 136Xe , using a Frequentist approach, as shown in Fig. 2.
 The likelihood is constructed as
@@
 The corresponding and are the bin values of the normalized energy spectrum convolved with the five-parameter energy response model.
-The Gaussian penalty term of the energy response contains the five-parameter and the covariant matrix in Run-.
-The Gaussian penalty terms are used to constrain the nuisance parameters , , and , which represent the relative uncertainties in the overall efficiency, the signal selection, and background model (see Table 2), respectively.
-The radioactivities of external structures, 124Xe concentrations, and solar and 7Be fluxes are identical in Run0 and Run1.
+The Gaussian penalty term of the energy response contains the five-parameter and the covariance matrix in Run-.
+The Gaussian penalty terms are used to constrain the nuisance parameters , , and , which represent the relative deviations in the overall efficiency, the signal selection, and the background model (see Table 2), respectively.
+The radioactivities of detector and SSP materials, 124Xe concentrations, and solar and 7Be neutrino fluxes are identical in Run0 and Run1.
 Other background components are treated independently for Run0 and Run1.
@@
 The horizontal axis represents the reconstructed energy in the data.
-Xe∗ includes the contributions from 124Xe, 125Xe, 127Xe, 129mXe, 131mXe, and 133Xe.
-The lower panel shows the residuals together with , and bands.
+Xe∗ denotes the contributions from 124Xe, 125Xe, 127Xe, 129mXe, 131mXe, and 133Xe.
+Material includes the contributions from SSP 232Th, SSP 238U, detector 60Co, detector 40K, detector 232Th, and detector 238U.
+The lower panel shows the residuals together with , , and bands.
 The observed backgrounds are in general agreement with expectations, as shown in Table 2.
 The fitted contributions from Kr in Run1 and SSP 232Th are pulled by 1.8 and 1.2, respectively, relative to their input values.
-The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
+The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), and the bias of the signal selection is (Run0) and (Run1).
 The half-life of 136Xe is measured as years, representing the most precise result so far, with a total uncertainty smaller than that of any previous measurement.
-It is worth noting that this result is slightly smaller than our previous result from Run0 data [PandaX:2022kwg], due to the use of the updated and slightly smaller 136Xe abundance of [PandaX:SSP], compared to the previous value of .
+It is worth noting that this result is slightly smaller than our previous result from Run0 data [40], due to the use of the updated and slightly smaller 136Xe abundance of [31], compared to the previous value of .
 The 136Xe spectrum is treated as the background in the search for Majoron-emitting signals, as shown later.
+In the fit of , we allow both the half-life of 136Xe and (i.e., the shape of 136Xe ) to float freely.
+The likelihood is almost the same as Eq. (3), but with a different definition of in Run-, which is modified as
+where and denote the normalized energy spectrum values in the bin for the first and second terms of the Taylor expansion of the decay rate, respectively.
+The parameter is defined as , where and denote the total detection efficiencies for the theoretical spectra of the first and second terms of the Taylor expansion, respectively.
+The rest of the parameters remain the same as in Eq. (4).
+The backgrounds, also summarized in Table 2, are consistent with expectations, except for the contributions from Kr in Run1 and SSP 232Th, which have pulls of 1.6 and 1.2, respectively.
+The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), and the bias of the signal selection is (Run0) and (Run1).
+We obtained a best fit of .
+Figure 3 compares the measured values of and the 136Xe half-life obtained by KamLAND-Zen [22] and by this work.
+Although both results are consistent with , our measured value has the opposite sign to that reported by KamLAND-Zen ().
+Assuming fully uncorrelated uncertainties, the two measurements differ by approximately .
+Our result is consistent with quasiparticle random-phase approximation predictions [22] at the level and remains compatible with shell-model calculations [22] within .
+However, the current precision does not allow us to exclude conclusively and prevents us from discriminating between the higher-state dominance and single-state dominance hypotheses [41].
+Achieving such discrimination will require improved precision in future experiments.
 Figure 3:
-Two-dimensional profile likelihood scan for 136Xe rate and .
-The black dot indicates the best-fit value of .
-The contours correspond to the 68.3, 90, 95.4, and 99.7 CL.
-In the fit of , we allow both the half-life of 136Xe and (i.e., the shape of 136Xe ) to float freely.
-The likelihood is almost the same as Eq. 3, but with a different definition of in Run- which is modified as
-where corresponds to the PDF corresponding to the first term of the Taylor expansion of the decay rate of , and corresponds to the PDF corresponding to the second term.
-The parameter , which comprises , and in Eq. 2 as well as the total detection efficiencies, varies linearly with .
-The rest of the parameters remain the same as in Eq. 4.
-The backgrounds, also summarized in Table 2, are consistent with expectations, except for the contributions from Kr in Run1 and SSP 232Th, which have pulls of 1.6 and 1.2, respectively.
-The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
-We obtained a best fit of , consistent with theoretical predictions [kesi31:theory] and the KamLAND-Zen results [KamLAND-Zen:kesi31].
-Fig. 3 shows the two-dimensional confidence intervals for the 136Xe rate and , revealing only a slight positive correlation.
-The inclusion of the second-order contribution is found to have a negligible impact on the 136Xe decay rate.
-The 136Xe decay half-life is estimated to be years.
-This result is consistent with the 136Xe -only fit, when the second-order contribution was not considered.
-Figure 4: Spectra fits with Majoron-emitting decay signals ().
-The figure is the same as Fig. 2 but with the upper limits (90% CL) of signals illustrated as hatched histograms. The lower panel shows the residuals together with , and bands.
-In the fit of Majoron-emitting , the likelihood is similar to Eq. 3, except that 136Xe is moved from the signal term to the background term, and then Majoron-emitting is introduced as the signal.
+Comparison of 136Xe half-lives and values from the KamLAND-Zen experiment [22] and our present analysis.
+The gray dashed line indicates the zero .
+The inclusion of the second-order contribution is found to have a negligible impact on the extracted 136Xe decay rate.
+The corresponding half-life is determined to be years, which is consistent with the result obtained from a 136Xe -only fit that neglects the second-order contribution.
+Figure 4: Spectra fits with Majoron-emitting decay signals () for Run0 (top) and Run1 (bottom), respectively.
+The figures are similar to Fig.
+2 but with the upper limit (90% CL) on the Majoron-emitting signal (mode ) illustrated as hatched histograms.
+The lower panel shows the residuals together with , , and bands.
+In the fit of Majoron-emitting , the likelihood is similar to Eq. (3), except that 136Xe is moved from the signal term to the background term, and then, Majoron-emitting is introduced as the signal.
 As an example, the best-fit spectrum for mode is shown in Fig. 4.
 No statistically significant evidence for Majoron-emitting decays is observed for any mode considered here, and the lower limits on Majoron-emitting half-lives are derived at the confidence level (CL), as summarized in Table 3.
-The extension of the lower bound of the ROI from 600 keV (as in EXO-200 [EXO-200:Majoron2021]) to 20 keV considerably improves the sensitivity to mode , where the Majoron-emitting spectrum peaks at approximately 600 keV, leading to the most stringent constraint.
-However, our sensitivity to modes 1, 2, and 3 is limited due to our 136Xe exposure being only 0.16 times that of EXO-200 [EXO-200:Majoron2021] and our higher background level.
+The extension of the lower bound of the ROI from 600 keV (as in EXO-200 [6]) to 20 keV considerably improves the sensitivity to mode , where the Majoron-emitting spectrum peaks at approximately 600 keV, leading to the most stringent constraint.
+However, our sensitivity to modes 1, 2, and 3 is limited due to our 136Xe exposure being only 0.16 times that of EXO-200 [6] and our higher background level.
 Similarly, the backgrounds are consistent with expectations, with only the contributions from Kr in Run1 and SSP 232Th pulled by 1.7 and 1.2, respectively (Table 2).
-The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), as well as the bias of the signal selection is (Run0) and (Run1).
-Table 3: The 90% CL lower limits on half-lives for different Majoron-emitting decay models and the comparison to EXO-200 results [EXO-200:Majoron2021]
+The bias of the overall efficiency obtained from the fit is (Run0) and (Run1), and the bias of the signal selection is (Run0) and (Run1).
+Table 3: The 90% CL lower limits on half-lives for different Majoron-emitting decay models and the comparison with EXO-200 results [6]
 Decay mode
@@
 EXO-200 (2021) (yr)
-n = 1
-n = 2
-n = 3
-n = 7
+= 1
+= 2
+= 3
+= 7
 In summary, we perform a measurement of in and a search for Majoron-emitting for 136Xe with a total 136Xe exposure of kgyr using PandaX-4T data.
 The most precise measurement of the 136Xe half-life to date is achieved thanks to the nearly complete spectrum used in the fit.
-Our measurement of is consistent with theoretical predictions and the results from KamLAND-Zen.
+Our measurement of is consistent with theoretical predictions.
+The values from KamLAND-Zen and this work are broadly consistent, with a modest (1.7) difference.
 The most competitive results for the Majoron-emitting mode with are obtained.
@@
 The capability to reconstruct the nearly complete 136Xe spectrum will allow us to fully exploit the PandaX-4T dataset, significantly enhancing its physics potential.
-This project is supported in part by grants from National Key R&D Program of China (Nos.
-2023YFA1606200, 2023YFA1606202), National Science Foundation of China (Nos.
-12090060, 12090062, 12305121, U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos.
-21TQ1400218, 22JC1410100, 23JC1410200, ZJ2023-ZD-003).
-We thank for the support by the Fundamental Research Funds for the Central Universities.
-We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund.
+Acknowledgements—This project is supported in part by grants from the National Key R&D Program of China (No.
+2023YFA1606200 and No.
+2023YFA1606202), National Science Foundation of China (No.
+12090060, No.
+12090062, No.
+12305121, No.
+U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (Grants No.
+21TQ1400218, No.
+22JC1410100, No.
+23JC1410200, No.
+ZJ2023-ZD-003).
+We are grateful for the support of the Fundamental Research Funds for the Central Universities.
+We also thank for their sponsorship: the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund.
 Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd.
 for indispensable logistical support and other help.
+Data availability—The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project.
+The data are available from the authors upon reasonable request.